OpenStudy (anonymous):

Prove the Converse of the Pythagorean Theorem using similar triangles. The Converse of the Pythagorean Theorem states that when the sum of the squares of the lengths of the legs of the triangle equals the squared length of the hypotenuse, the triangle is a right triangle. Be sure to create and name the appropriate geometric figures.

OpenStudy (anonymous):

@ash2326 @aaronq @andriod09 @BulletWithButterflyWings @BostonBlue @e.mccormick @Erin001001 @stabar @iiamentertainment @jhonyy9 @goformit100 @hba @Haunted9 @lncognlto @rizwan_uet can someone help

OpenStudy (anonymous):

i am not good at math so idk

OpenStudy (anonymous):

ok thx anyways :)

OpenStudy (zzr0ck3r):

I hate to just post that, but I cant draw on my tablet

OpenStudy (zzr0ck3r):

tell me if you dont understand something

OpenStudy (anonymous):

ok but thank you :)

OpenStudy (anonymous):

EF = BC = a ÐF is a right angle. FD = CA = b triangle EF = BC = a angle F is a right angle. FD = CA = b In triangle DEF, By Pythagoras Theorem, a2 + b2 = c2 the given AB=c= a^2 + b^2 square root Theorefore AB = DE But by construction, BC = EF and CA = FD triangle ABC congruent to DEF (S.S.S.)

OpenStudy (anonymous):

like this?

OpenStudy (zzr0ck3r):

yep, as long as you understand it.

OpenStudy (anonymous):

and i have this question

OpenStudy (anonymous):

i do and thanks to you :D the link explains everything :)

OpenStudy (anonymous):

@skullpatrol @phi @Flisk

OpenStudy (phi):

use asa

OpenStudy (anonymous):

can you show and explain ? :)

OpenStudy (phi):

you are given angle B= < F <DCE = < DEC any guesses what angle-side-angle we are going to use ?

OpenStudy (anonymous):

prove two angles are the same and a side?

OpenStudy (phi):

yes, but what angles and side are we going to use ?

OpenStudy (phi):

<B= < F is given. so we should use those angles they tell us <DCE = < DEC those angles are "vertical angles" to an angle in the triangle we are trying to show are congruent. that looks promising. so it seems we want to use the side between <B and <ACB and the corresponding side in the other triangle

OpenStudy (anonymous):

so thats given to

OpenStudy (anonymous):

ok :)

OpenStudy (phi):

almost. DF = DB, but we need to show EF= CB

OpenStudy (phi):

but notice EF= DF - DC and CB= DB-DE if you know DC= DE, and DF= DB then the difference will be the same

OpenStudy (phi):

so you need to know why DC= DE I would use isosceles triangle idea... if base angles are equal, then the opposite sides are equal

OpenStudy (anonymous):

now i have to put the proof in order

OpenStudy (phi):

yes, lots of details but the big picture is angle - side -angle

OpenStudy (anonymous):

given: angle B= < F <B= < F is given this is what i wrote til now

OpenStudy (anonymous):

<DCE = < DEC are verical angles EF= DF - DC and CB= DB-DE DC= DE opposite sides are equal triangle ABE congruent to triangle GFE by (ASA)

OpenStudy (phi):

I hope you mean <DCE = < DEC given then you say <ACB = <DCE vertical angles also <FEG= < DEC vertical angles now you want to say <ACB = <FEG why is this true ?

OpenStudy (phi):

we are doing the angles first. why is angle ACB = angle FEG ? you start by saying <DCE = < DEC given <ACB = <DCE vertical angles <FEG = < DEC vertical angles

OpenStudy (phi):

if <DCE = <DEC you could say <ACB= <DCE means <ACB= <DEC things equal to the same thing (in this case both <DEC and <ACB equal <DEC) are equal to each other so <ACB= <DEC now you have <ACB= <DEC things equal to the same thing are equal to each other <FEG = < DEC vertical angles now why is <ACB = <FEG ?

OpenStudy (anonymous):

angle sides angle?

OpenStudy (phi):

I guess this does not make sense to you. notice that <ACB = <DEC and <FEG = *the same thing* < DEC do you agree <ACB and <FEG both equal the same thing <DEC ? why is <ACB = <FEG ?

OpenStudy (anonymous):

well yah your rite it does not make sense but i get the first part up to vertical angles

OpenStudy (anonymous):

really ok that looks simple !

OpenStudy (phi):

you need the idea that if two different things equal the same thing then they equal each other. for example, if A = C and B=C then A=B this is common sense, so it should make a little sense. For this problem you say <DCE = < DEC given <ACB= <DCE vertical angles <ACB= <DEC things equal to the same thing are equal to each other <FEG = < DEC vertical angles <ACB = <FEG things equal to the same thing are equal to each other

OpenStudy (phi):

that is the whole sequence to show <ACB=<FEG

OpenStudy (phi):

now tackle the side part. First step. <DCE = < DEC given side DC= side DE why ?

OpenStudy (anonymous):

but isnt there at the end ASA?

OpenStudy (anonymous):

vertical angles

OpenStudy (phi):

the stuff to show <ACB=<FEG is part I. the first angle. now we need part II, the side

OpenStudy (anonymous):

corresponding sides so..

OpenStudy (phi):

First step. <DCE = < DEC given side DC= side DE why ? sides are not vertical angles. see http://www.regentsprep.org/Regents/math/geometry/GP6/Lisosceles.htm and look carefully for "converse theorem"

OpenStudy (anonymous):

If two angles of a triangle are congruent, the sides opposite them are congruent. this

OpenStudy (phi):

yes. Do you follow what it says? we are given the base angles are =, so using this theorem (which has been proven by some smart person) we *know* the opposite sides (DC and DE) are equal

OpenStudy (anonymous):

yahh i get it now :) thank you for that! ^_^

OpenStudy (phi):

now we know DC +CB= DB

OpenStudy (anonymous):

ok

OpenStudy (phi):

now we know DC +CB= DB sum of parts equal the whole and DE+EF= DF sum of parts equal the whole DB= DF given so we can say DC +CB= DF (replace DB with DF) and DC+CB = DE+EF things equal to the same thing are equal to each other we just proved DC= DE so we can say DC+CB= DC+EF (replace DE with DC) CB= EF (subtracting equal amounts keeps the equality)

OpenStudy (phi):

that proves the sides the last part, is <B= <F given so you have shown angle-side-angle are congruent, so the two triangles ABC= GFE

OpenStudy (anonymous):

wow that was a long answer and im sure your tired of all this typing im sorry and thank you so so so much for all you help!!

OpenStudy (phi):

yes, lots of details.

OpenStudy (anonymous):

@bayan are u still in geometry can u help me